The measurement of the smallest angle in a trianle is 50 degrees less than the measure of he largest angle. The measure of the remaining angle is 10 degrees more than the smallest angle. What is the measurement od each angle?
Answer by rfer(16322)

x-50
(x-50)+10
x
----
x-50+x-50+10+x=180
3x=270
x=90 degrees
x-50=40 degrees
(x-50)+10=50 degrees